Considering the following hypotheses:
H0: μ = 73
HA: μ ≠ 73
What are the p-values based on the following sample information? (Feel free to reference the appropriate table: z table or t table)

1. x¯ = 70; s = 6.4; n = 19
2. x¯ = 70; s = 6.4; n = 24
3. x¯= 70; s = 6.9; n = 35
4. x¯ = 76; s = 6.9; n = 35

Answer by Boreal(15235)

with s use a t-distribution
t (df=18)=-3/6.4/sqrt(19)=-2.04
p=0.056
with n=24, the t=-2.30 and
p=0.0311
with the third
t=-3/6.9/sqrt(35)=-2.57
p-value is 0.0146
The fourth is the same, since the t-distribution is symmetric and the difference from the mean is +3 now, rather than -2.